Write a quadric equation having the given solutions:
-2,13

Answer by RAY100(1637)

x=-2,,,(x+2) =0
.
x=13,,,,(x-13) =0
.
(x+2)(x-13) =0
.
x^2 -11x -26 =0
.
checking
.
for (-2) ,,,-2^2 -11(-2) -26 =0 ,,,,4+22-26=0,,,,,ok
.
for (13),,,,(13)^2 -11(13) -26 =0,,, 169-143 -26 =0,,, 0=0,,,ok